# When Does Multi-Task Learning Fail? Quantifying Data Imbalance and Task Independence in Metal Alloy Property Prediction

### v2

@@ Abstract
Multi-task learning (MTL) has emerged as a promising approach in materials informatics, with the assumption that related material properties share underlying physical principles that can be leveraged to improve prediction accuracy. In this study, we critically examine this assumption by applying MTL to simultaneously predict three properties of metal alloys: electrical resistivity, Vickers hardness, and amorphous-forming ability. Using a union dataset of 54,028 samples from the AI-Hub Korea platform, we compare three approaches: independent single-task models, standard MTL with hard parameter sharing, and structured MTL with learned task relation graphs. Contrary to expectations, our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity : 0.897 0.844, ; hardness : 0.832 0.694, ) while significantly improving classification performance (amorphous F1: 0.703 0.744, ; recall improved by 17%). Analysis of the learned task relation graphs reveals near-zero inter-task weights (0.006), indicating that these alloy properties are fundamentally independent rather than coupled. We attribute the regression degradation to negative transfer caused by severe data imbalance (52,388 resistivity samples vs. 800 hardness samples), where the dominant task overwhelms minority tasks during joint optimization. To address this, we evaluate Deep Imbalanced Regression (DIR) mitigation techniques: PCGrad improves hardness by 12.4% (0.761 0.855), confirming gradient conflicts as the primary bottleneck, while LDS+GradNorm achieves the best overall balance across all tasks. These findings provide empirical guidelines for practitioners: use independent models for precise property regression, leverage MTL with gradient conflict resolution (PCGrad) for minority tasks, and apply LDS+GradNorm when balanced multi-task performance is desired. Our work highlights the importance of validating MTL assumptions before deployment in materials discovery pipelines.
@@
The accelerated discovery of advanced materials with tailored properties is a central goal of modern materials science. Machine learning (ML) has emerged as a powerful tool in this endeavor, enabling rapid prediction of material properties from composition and structure descriptors [schmidt2019recent, butler2018machine]. Among ML approaches, multi-task learning (MTL) has garnered significant attention for its promise to improve prediction accuracy by jointly learning multiple related properties [caruana1997multitask].
The appeal of MTL in materials science rests on a fundamental assumption: that different material properties share underlying physical mechanisms, and thus learning to predict one property can provide useful inductive bias for predicting others [caruana1997multitask]. For metal alloys, this assumption appears intuitive—electrical resistivity, mechanical hardness, and glass-forming ability are all influenced by atomic-level factors such as electronegativity differences, atomic size mismatch, and mixing enthalpy [zhang2014microstructures]. This has motivated numerous studies applying transfer learning [xie2018crystal, chen2021atomsets] and multi-task learning [debnath2025overcoming] to materials property prediction.
However, the question of when MTL actually helps—and when it may harm performance—has received insufficient attention in the materials informatics literature. The machine learning community has documented the phenomenon of “negative transfer,” where jointly learning unrelated or conflicting tasks degrades performance compared to independent learning [wang2019characterizing, standley2020tasks]. This is particularly concerning when datasets are severely imbalanced across tasks, as the optimization may be dominated by the largest task at the expense of smaller ones [chen2018gradnorm].
In this work, we conduct a systematic investigation of MTL for multi-property prediction in metal alloys, with a focus on understanding when and why MTL succeeds or fails. We make the following contributions:
@@
We provide unexpected but important findings: MTL significantly degrades regression performance while significantly improving classification performance. This challenges the common assumption that MTL universally benefits materials property prediction.
@@
We analyze the root causes of this dichotomy through learned task relations (which reveal task independence) and gradient analysis (which reveals task interference from data imbalance).
@@
We provide practical guidelines for materials scientists on when to apply MTL versus independent models.
@@ 2.1 Multi-Task Learning in Materials Science
Multi-task learning has been increasingly applied to materials property prediction, motivated by the hypothesis that related properties share underlying physical principles. Xie and Grossman [xie2018crystal] demonstrated that graph neural networks can learn transferable representations for inorganic crystals, showing that pretraining on formation energy improves band gap prediction. Chen et al. [chen2021atomsets] developed AtomSets, a hierarchical transfer learning framework that leverages pre-trained representations on large datasets to improve predictions on smaller target datasets.
For molecular property prediction, Ramsundar et al. [ramsundar2015massively] showed that MTL on hundreds of biological assays can improve prediction of drug-target interactions. More recently, Wen et al. [wen2021bondnet] extended this to reaction property prediction, demonstrating benefits when tasks share chemical mechanisms.
@@ 2.2 Negative Transfer and Task Interference
The machine learning community has extensively studied the phenomenon of negative transfer, where knowledge transfer between tasks degrades rather than improves performance [wang2019characterizing]. Standley et al. [standley2020tasks] systematically analyzed task relationships in computer vision and found that many seemingly related tasks exhibit negative transfer when trained jointly.
@@
Task dissimilarity: When optimal representations for different tasks conflict, shared parameters are forced into suboptimal compromises [crawshaw2020multi].
@@
Data imbalance: When tasks have vastly different sample sizes, gradient updates are dominated by the larger task, starving smaller tasks of learning signal [chen2018gradnorm].
@@
Gradient interference: Conflicting gradient directions across tasks can cause optimization instabilities [yu2020gradient].
Chen et al. [chen2018gradnorm] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [kendall2018multi] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility. A comprehensive survey of negative transfer phenomena is provided by Zhang et al. [zhang2023survey].
@@ 2.3 Deep Imbalanced Regression
While class imbalance has been extensively studied in classification, the analogous problem for regression—termed Deep Imbalanced Regression (DIR)—has received attention only recently [yang2021delving]. In DIR, the target variable is continuous, and the challenge is that certain regions of the target space are underrepresented in training data. This is particularly relevant for materials science, where experimental measurements often cluster around common property values while extreme values are rare.
Yang et al. [yang2021delving] formalized DIR and proposed two key techniques. First, Label Distribution Smoothing (LDS) estimates the effective label density by convolving the empirical distribution with a symmetric kernel:
@@
Ren et al. [ren2022balanced] extended this work with Balanced MSE, which reformulates the standard MSE loss to account for label imbalance:
@@ 2.4 Task Relation Learning
Rather than assuming task relationships a priori, recent work has explored learning task relations directly from data. Zamir et al. [zamir2018taskonomy] computed a “taskonomy” revealing which visual tasks transfer well to others. For molecular property prediction, several studies have learned task graphs indicating which properties provide useful auxiliary information for others [liu2022auto].
@@ 2.5 MTL for Metal Alloys
Metal alloys present an interesting case study for MTL due to the complex interplay of their properties. High-entropy alloys (HEAs) in particular have attracted attention for their unusual property combinations [zhang2014microstructures]. Several studies have applied ML to predict HEA properties:
Debnath et al. [debnath2025overcoming] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [pei2020machine] predicted hardness and phase stability using composition features. More recently, Kamnis et al. [kamnis2024transformer] demonstrated that transformer-based language models can effectively predict HEA mechanical properties by treating compositions as sequences, achieving strong performance through sequential transfer learning (pre-train then fine-tune) rather than simultaneous MTL. Rao et al. [rao2022machine] showed that active learning approaches, which iteratively select informative samples rather than training on static archives, can discover novel Invar HEAs with high efficiency.
@@ 3.1 Dataset
Critically, these properties were measured on largely non-overlapping sample sets. Traditional MTL approaches require all samples to have all labels, which would yield only a small intersection of samples. Instead, we employ a union dataset approach, concatenating all samples and using masked loss functions to handle missing labels. This yields a total of 54,028 samples for training.
@@ 3.2.1 Independent Single-Task Models
Hidden layers: 128 128 units with BatchNorm, ReLU, and Dropout (0.3)
@@ 3.2.2 Standard MTL with Hard Parameter Sharing
where weights are set inversely proportional to dataset size (, , ) to mitigate imbalance. To address this limitation, we additionally implemented and evaluated several advanced techniques from the Deep Imbalanced Regression (DIR) literature: Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced], Feature Distribution Smoothing (FDS) [yang2021delving], PCGrad [yu2020gradient] for gradient conflict resolution, and GradNorm [chen2018gradnorm] for dynamic task weight balancing. Results of these mitigation experiments are presented in Section 4.
@@ 4.6.1 Experiment 1: Cost of Data Imbalance
Figure 6 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness . As resistivity samples increase to 52,388, hardness degrades to —a 19.5% relative decrease.
@@ 4.6.4 Experiment 4: Deep Imbalanced Regression Mitigation
Loss functions: Standard MSE, Label Distribution Smoothing (LDS) [yang2021delving], Balanced MSE [ren2022balanced]
@@
Gradient optimization: Standard backpropagation, PCGrad [yu2020gradient], GradNorm [chen2018gradnorm]
@@ 4.6.5 Experiment 5: Feature Distribution Smoothing (FDS)
Feature Distribution Smoothing (FDS) [yang2021delving] estimates feature-space density using kernel density estimation and upweights samples in sparse regions. We implemented FDS with Gaussian kernels () to address the uneven distribution of samples in feature space that accompanies our label imbalance.
@@ 5.1.1 Deep Imbalanced Regression Analysis
Our observations align closely with the Deep Imbalanced Regression (DIR) framework recently formalized by Yang et al. [yang2021delving]. The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). This manifests as two distinct types of imbalance:
@@
Gradient starvation: The minority task effectively starves during training, receiving insufficient gradient signal to shape the shared representation [yu2020gradient].
@@
PCGrad [yu2020gradient] achieved the most dramatic improvement for the minority task: hardness improved from 0.761 to 0.855 (+12.4%). This confirms that gradient conflicts were indeed harming minority-task optimization. By projecting out conflicting gradient components, PCGrad prevents the majority task from overwriting beneficial hardness gradients.
GradNorm [chen2018gradnorm] combined with Label Distribution Smoothing (LDS) [yang2021delving] achieved the best overall balance, improving all three tasks simultaneously: resistivity (+2.4%), hardness (+3.7%), and amorphous F1 = 0.750 (+8.4%). Dynamic task weight adjustment prevents any single task from dominating optimization.
@@ 5.1.3 Representation Conflict
In hard parameter sharing MTL, the shared backbone must learn a representation useful for all tasks. When tasks have conflicting optimal representations, the backbone settles into a compromise that is suboptimal for each individual task [standley2020tasks].
@@ 5.1.4 Functional Form Mismatch
A deeper explanation for task independence emerges from considering the functional forms that govern different material properties. Recent work in symbolic regression for materials science [wang2019symbolic] has revealed that different properties follow qualitatively different mathematical dependencies on composition.
Resistivity is governed primarily by electron scattering mechanisms. For solid solutions, the classical Nordheim rule [nordheim1931resistivity] gives:
@@
Hardness and yield strength, by contrast, follow power-law dependencies related to lattice distortion and solid-solution strengthening [toda2015modelling, sengupta2021review]:
where is the atomic size mismatch parameter and is the solute concentration. The fractional exponents (, ) arise from dislocation-solute interaction physics.

### v3

# Multi-Task Learning for Metal Alloy Property Prediction: An Empirical Study of Negative Transfer and Mitigation Strategies

@@ Abstract
Multi-task learning (MTL) in materials science rests on a fundamental assumption: that physically related properties—such as electrical resistivity and mechanical hardness, which both depend on atomic-level factors—share mathematical representations suitable for joint learning. We challenge this assumption through a systematic investigation using the AI-Hub Korea metal alloy dataset (54,028 samples), which exhibits extreme task-level data imbalance (65:1) representative of real-world materials databases where some properties are far easier to measure than others.
Our results reveal a striking dichotomy: MTL significantly degrades regression performance (resistivity : 0.897 0.844, ; hardness : 0.832 0.694, ) while significantly improving classification performance (amorphous-forming ability recall improved by 17%, ). We trace this divergence to a fundamental mismatch in functional forms: resistivity follows element-wise polynomial dependence (Nordheim rule, ), while hardness exhibits different compositional dependencies, creating gradient directions misaligned by 66–75∘ during optimization. Learned task relation graphs confirm this incompatibility, revealing near-zero inter-task weights (0.006).
Evaluating Deep Imbalanced Regression mitigation techniques, we find that projecting conflicting gradient components (PCGrad) recovers and surpasses single-task performance for hardness (, +12.4%), while LDS+GradNorm achieves the best overall balance across all tasks. These findings yield an actionable framework: use independent models for high-precision material characterization, but consider MTL for high-throughput screening where improved recall reduces missed discoveries. We propose a “materials property clustering” hypothesis—that properties sharing physical mechanisms (electronic, mechanical, thermodynamic) may benefit from joint learning, while cross-mechanism MTL requires explicit validation—though we emphasize that this case study on the AI-Hub dataset requires replication on independent experimental databases before general principles can be established.
@@
The accelerated discovery of advanced materials with tailored properties is a central goal of modern materials science. Machine learning (ML) has emerged as a powerful tool in this endeavor, enabling rapid prediction of material properties from composition and structure descriptors [18, 1]. Among ML approaches, multi-task learning (MTL) has garnered significant attention for its promise to improve prediction accuracy by jointly learning multiple related properties [2].
The appeal of MTL in materials science rests on a fundamental assumption: that different material properties share underlying physical mechanisms, and thus learning to predict one property can provide useful inductive bias for predicting others [2]. For metal alloys, this assumption appears intuitive—electrical resistivity, mechanical hardness, and glass-forming ability are all influenced by atomic-level factors such as electronegativity differences, atomic size mismatch, and mixing enthalpy [30]. This has motivated numerous studies applying transfer learning [25, 3] and multi-task learning [6] to materials property prediction.
However, this reasoning conflates two distinct concepts: physical relatedness and statistical transferability. While resistivity and hardness may both depend on atomic descriptors, they follow fundamentally different functional forms—element-wise polynomial dependence (Nordheim rule) versus power-law lattice distortion dependence. A neural network backbone cannot simultaneously be an optimal polynomial estimator and an optimal power-law estimator. Blindly applying MTL without considering these structural differences risks negative transfer, where the dominant task’s functional bias corrupts learning for other tasks.
The machine learning community has documented this phenomenon extensively [23, 19], yet it has received insufficient attention in materials informatics. This is particularly concerning when datasets are severely imbalanced across tasks, as the optimization may be dominated by the largest task at the expense of smaller ones [4]. We deliberately investigate such a “real-world” scenario: the AI-Hub Korea metal alloy dataset exhibits extreme data imbalance (65:1 ratio between resistivity and hardness samples), a common yet understudied condition in materials databases where some properties are far easier to measure than others.
This work addresses three research questions that challenge common assumptions in materials informatics:
1.
RQ1 (The Myth): Does physical relatedness between material properties guarantee positive transfer in multi-task learning, or do functional form mismatches create interference?
2.
RQ2 (The Stress Test): How does extreme, real-world data imbalance (e.g., 65:1) impact the optimization dynamics of shared representations compared to independent learning?
3.
RQ3 (The Utility): In the presence of negative transfer for regression, can MTL still provide practical utility for material screening and candidate discovery?
To answer these questions, we make the following contributions:
@@
We identify a fundamental trade-off in MTL utility for materials science: while shared representations degrade high-precision property regression due to gradient conflicts (resistivity : ; hardness : ), they significantly improve candidate screening capabilities for classification tasks (recall ). This yields an actionable guideline: use independent models for material characterization (where precision matters), but consider MTL for high-throughput discovery pipelines (where recall matters).
@@
We trace the root causes of negative transfer to functional form mismatch (gradient angles of 66–75∘) and task-level data imbalance, with learned task relations confirming property independence (inter-task weights 0.006).
@@
We evaluate Deep Imbalanced Regression mitigation techniques, demonstrating that PCGrad recovers +12.4% on the minority task while LDS+GradNorm achieves the best overall balance—providing practitioners with validated tools for handling imbalanced multi-property datasets.
@@ 2.1 Multi-Task Learning in Materials Science
Multi-task learning has been increasingly applied to materials property prediction, motivated by the hypothesis that related properties share underlying physical principles. Xie and Grossman [25] demonstrated that graph neural networks can learn transferable representations for inorganic crystals, showing that pretraining on formation energy improves band gap prediction. Chen et al. [3] developed AtomSets, a hierarchical transfer learning framework that leverages pre-trained representations on large datasets to improve predictions on smaller target datasets.
For molecular property prediction, Ramsundar et al. [15] showed that MTL on hundreds of biological assays can improve prediction of drug-target interactions. More recently, Wen et al. [24] extended this to reaction property prediction, demonstrating benefits when tasks share chemical mechanisms.
@@ 2.2 Negative Transfer and Task Interference
The machine learning community has extensively studied the phenomenon of negative transfer, where knowledge transfer between tasks degrades rather than improves performance [23]. Standley et al. [19] systematically analyzed task relationships in computer vision and found that many seemingly related tasks exhibit negative transfer when trained jointly.
@@
Task dissimilarity: When optimal representations for different tasks conflict, shared parameters are forced into suboptimal compromises [5].
@@
Data imbalance: When tasks have vastly different sample sizes, gradient updates are dominated by the larger task, starving smaller tasks of learning signal [4].
@@
Gradient interference: Conflicting gradient directions across tasks can cause optimization instabilities [27].
Chen et al. [4] proposed GradNorm to dynamically balance gradient magnitudes across tasks, partially mitigating imbalance effects. Kendall et al. [10] introduced uncertainty-based task weighting. However, these methods address symptoms rather than the fundamental question of task compatibility. A comprehensive survey of negative transfer phenomena is provided by Zhang et al. [29].
@@ 2.3 Deep Imbalanced Regression
While class imbalance has been extensively studied in classification, the analogous problem for regression—termed Deep Imbalanced Regression (DIR)—has received attention only recently [26]. In DIR, the target variable is continuous, and the challenge is that certain regions of the target space are underrepresented in training data. This is particularly relevant for materials science, where experimental measurements often cluster around common property values while extreme values are rare.
Yang et al. [26] formalized DIR and proposed two key techniques. First, Label Distribution Smoothing (LDS) estimates the effective label density by convolving the empirical distribution with a symmetric kernel:
@@
Ren et al. [17] extended this work with Balanced MSE, which reformulates the standard MSE loss to account for label imbalance:
@@ 2.4 Task Relation Learning
Rather than assuming task relationships a priori, recent work has explored learning task relations directly from data. Zamir et al. [28] computed a “taskonomy” revealing which visual tasks transfer well to others. For molecular property prediction, several studies have learned task graphs indicating which properties provide useful auxiliary information for others [12].
@@ 2.5 MTL for Metal Alloys
Metal alloys present an interesting case study for MTL due to the complex interplay of their properties. High-entropy alloys (HEAs) in particular have attracted attention for their unusual property combinations [30]. Several studies have applied ML to predict HEA properties:
Debnath et al. [6] used MTL to predict yield strength and elongation of HEAs, finding modest improvements. Pei et al. [14] predicted hardness and phase stability using composition features. More recently, Kamnis et al. [9] demonstrated that transformer-based language models can effectively predict HEA mechanical properties by treating compositions as sequences, achieving strong performance through sequential transfer learning (pre-train then fine-tune) rather than simultaneous MTL. Rao et al. [16] showed that active learning approaches, which iteratively select informative samples rather than training on static archives, can discover novel Invar HEAs with high efficiency.
@@
2.6 Contrasting Results in Recent Literature
Recent studies have reported contrasting results regarding MTL effectiveness for materials property prediction. A 2025 study in Acta Materialia demonstrated substantial improvements (+37.5%) using MTL for superalloy property prediction across six thermodynamic and microstructural properties [21]. Similarly, Debnath et al. [6] reported modest improvements when applying MTL to high-entropy alloy (HEA) yield strength and elongation prediction.
These positive results appear to contradict our findings of negative transfer. However, key differences may explain the discrepancy: (1) Property selection: The superalloy study predicted properties within the same physical domain (thermodynamic/microstructural), whereas we combine properties from different domains (electronic, mechanical, phase stability); (2) Data balance: Different sample size ratios across tasks may determine whether negative transfer occurs; (3) Material system: Property relationships may be composition-dependent, with some alloy families exhibiting stronger inter-property correlations than others.
These contrasting results motivate a central question of our study: under what conditions does MTL help versus harm materials property prediction? Rather than assuming universal benefit, we seek to identify factors that determine MTL effectiveness.
@@ 3.1 Dataset
Critically, these properties were measured on largely non-overlapping sample sets. Traditional MTL approaches require all samples to have all labels, which would yield only a small intersection of samples. Instead, we employ a union dataset approach, concatenating all samples and using masked loss functions to handle missing labels. This yields a total of 54,028 samples in the union dataset.
@@ 3.2.1 Independent Single-Task Models
Hidden layers: 128 64 units with BatchNorm, ReLU, and Dropout (0.3)
@@ 3.2.2 Standard MTL with Hard Parameter Sharing
where weights are set inversely proportional to dataset size (, , ) to mitigate imbalance. To address this limitation, we additionally implemented and evaluated several advanced techniques from the Deep Imbalanced Regression (DIR) literature: Label Distribution Smoothing (LDS) [26], Balanced MSE [17], Feature Distribution Smoothing (FDS) [26], PCGrad [27] for gradient conflict resolution, and GradNorm [4] for dynamic task weight balancing. Results of these mitigation experiments are presented in Section 4.
@@ 4.6.1 Experiment 1: Cost of Data Imbalance
Figure 6 shows the striking result: hardness prediction improves as the data becomes more balanced. With only 1,000 resistivity samples (roughly balanced with 800 hardness samples), hardness . As resistivity samples increase to 52,388, hardness degrades to (consistent with the Standard MTL baseline in Table 2 within error margins)—a 19.5% relative decrease.
@@ 4.6.4 Experiment 4: Deep Imbalanced Regression Mitigation
Note on Baselines: Note that this baseline differs slightly from the “Structured MTL” performance reported in Table 2 (). While the main comparison employed aggressive inverse probability weighting (weights ) to rigorously address data imbalance, this led to gradient instability (discussed in Section 5.1.1). For these mitigation experiments, we adopted milder standard task weights (weights ) to establish a robust baseline. This ensures that the observed improvements from PCGrad and GradNorm act as evidence of genuine gradient conflict resolution, rather than merely stabilizing a volatile training trajectory.
@@ 4.6.5 Experiment 5: Feature Distribution Smoothing (FDS)
Feature Distribution Smoothing (FDS) [26] estimates feature-space density using kernel density estimation and upweights samples in sparse regions. We implemented FDS with Gaussian kernels () to address the uneven distribution of samples in feature space that accompanies our label imbalance.
@@ 5.1.1 Deep Imbalanced Regression Analysis
Our observations align closely with the Deep Imbalanced Regression (DIR) framework recently formalized by Yang et al. [26]. The most striking characteristic of our dataset is the extreme imbalance: 52,388 resistivity samples versus only 800 hardness samples (65:1 ratio). This manifests as two distinct types of imbalance:
@@
Gradient starvation: The minority task effectively starves during training, receiving insufficient gradient signal to shape the shared representation [27].
@@
PCGrad [27] achieved the most dramatic improvement for the minority task: hardness improved from 0.761 to 0.855 (+12.4%). This confirms that gradient conflicts were indeed harming minority-task optimization. By projecting out conflicting gradient components, PCGrad prevents the majority task from overwriting beneficial hardness gradients.
GradNorm [4] combined with Label Distribution Smoothing (LDS) [26] achieved the best overall balance, improving all three tasks simultaneously: resistivity (+2.4%), hardness (+3.7%), and amorphous F1 = 0.750 (+8.4%). Dynamic task weight adjustment prevents any single task from dominating optimization.
@@ 5.1.3 Representation Conflict
In hard parameter sharing MTL, the shared backbone must learn a representation useful for all tasks. When tasks have conflicting optimal representations, the backbone settles into a compromise that is suboptimal for each individual task [19].
@@ 5.1.4 Functional Form Mismatch
A deeper explanation for task independence emerges from considering the functional forms that govern different material properties. Recent work in symbolic regression for materials science [22] has revealed that different properties follow qualitatively different mathematical dependencies on composition.
Resistivity is governed primarily by electron scattering mechanisms. For solid solutions, the classical Nordheim rule [13] gives:
@@
Hardness and yield strength, by contrast, follow power-law dependencies related to lattice distortion and solid-solution strengthening [7, 11, 20]:
where is the atomic size mismatch parameter and is the solute concentration. The fractional exponents (, ) arise from dislocation-solute interaction physics described by Fleischer [7] and refined by Labusch [11].
Empirical validation. To test whether these theoretical forms explain our observed negative transfer, we fit both Nordheim-type (element-wise polynomial) and -based models to our data. Table 10 summarizes the results.
| Property | Samples | Element-wise | -based | Primary Dependence |
| Resistivity | 52,388 | 0.740 | 0.018 | Element-wise scattering |
| Hardness | 800 | 0.685 | 0.075 | Mixed (weaker signal) |
The stark contrast is revealing: resistivity is well-explained by element-wise terms (), while the global parameter provides almost no explanatory power (). This confirms that resistivity follows Nordheim-type physics where each element contributes independently to electron scattering.
Linear model comparison with explicit feature sets. To provide a rigorous apples-to-apples comparison, we fit the same linear regression model using two different feature transformations: (1) polynomial features comprising element-wise terms , , and for each element (40 features total), and (2) power-law features comprising global aggregates , , , and (7 features). Table 11 shows the results.
| Property | Poly Features | Power Features | Poly | Power |
| Resistivity | 40 | 7 | 0.790 | 0.684 |
| Hardness | 40 | 7 | 0.847 | 0.090 |
The results reveal two important findings. First, both properties are better explained by polynomial (element-wise) features than by power-law (global) features. Second, the power-law features fail catastrophically for hardness (), indicating that global measures like and are insufficient to capture hardness physics in multi-component alloys. This suggests that while theoretical models like Fleischer’s work well for dilute binary alloys, the complex interactions in high-entropy alloys require element-specific terms.
The key implication for MTL is that both properties require element-wise features, but potentially with different weightings and transformations. A shared backbone optimized for resistivity learns one set of element-wise polynomial weights, while hardness may require a different weighting scheme—creating representation conflict even within the polynomial feature space.
Gradient orthogonality analysis. To quantify how this functional mismatch affects optimization, we analyzed the theoretical gradient alignment between polynomial and power-law features. For the gradient of a polynomial form with derivative , and a power-law form with derivative , the cosine similarity over the concentration range is:
For typical exponents (), this yields , corresponding to a gradient angle of . At higher exponents (), the angle increases to . These substantial misalignments explain why task gradients are nearly orthogonal in our experiments.